The difference between two numbers is 15. When twice the smaller number is subtracted from three times the larger number the difference is 58. What is the larger number?
Answer by alicealc(167)

1st number = x
2nd number = y

x - y = 15
3x - 2y = 58
multiply the 1st equation by 2:
2x - 2y = 30
then subtract it from the 2nd equation:
3x - 2y = 58
2x - 2y = 30
-------------- (-)
x = 28

substitute x = 28 into x - y = 15
28 - y = 15
28 - 15 = y
y = 13

so,the larger number is 28